# Finite-sample valid prediction of future insurance claims in the regression problem

### v1

@@ Abstract
In the current insurance literature, prediction of insurance claims in the regression problem is often performed with a statistical model. This model-based approach may suffer from several drawbacks: (i) model misspecification, (ii) selection effect, and (iii) lack of finite-sample validity. This article addresses these three issues simultaneously by employing conformal prediction—a general machine learning strategy for valid predictions. The proposed method is both model-free and tuning-parameter-free. It also guarantees finite-sample validity at a pre-assigned coverage probability level.
@@ 1 Introduction
Prediction is one of the most important inferential tasks for actuaries since it forms the basis for many key aspects of an insurer’s business operations, such as premium calculation and reserves estimation. According to Shmueli (2010), there are two key goals in data science and statistics: to explain and to predict. However, these two goals often warrant different approaches. For example, as demonstrated in Shmueli (2010), a wrong model, under some conditions, can even beat the oracle model in prediction, but the same cannot be said for explanation. This paper only concerns prediction. In the existing insurance literature, prediction is often performed using either a parametric approach or a non-parametric approach (e.g., Frees et al. 2014). In the parametric approach, the actuary posits a model, applies model selection tools to choose the “best” model, trains the chosen model, and finally makes predictions; see, for example, Claeskens and Hjort (2008) and Part I of Frees (2010). While this parametric approach has been widely applied in insurance, it has several drawbacks. First, the posited model may be misspecified, leading to grossly misleading predictions (Hong and Martin 2020). Second, scholars recently discovered that model selection might introduce serious biases in prediction; this adverse effect is called the selection effect (e.g., Berk et al. 2013; Kuchibhotla et al. 2022); see, for example, Hong and Martin (2018a, b) for some simple illustrations in insurance. Though several proposals have been advanced to alleviate the selection effect, a satisfying answer has yet to be found (Kuchibhotla et al. 2022). Third, prediction is uncertain and should be probabilistically calibrated to achieve finite-sample validity (Gneiting and Katzfuss 2014; Vovk 2024); unfortunately, no parametric model is known to achieve finite-sample validity. For these reasons, some actuaries opt for a non-parametric model for predicting insurance claims (e.g., Jeon and Kim 2013; Hong and Martin 2017; Richardson and Hartman 2018). While most non-parametric models shun both the model misspecification risk and the selection effect, they can still fall short of our expectations for two reasons. In the first place, many of these models contain some tuning parameters and the choices of these tuning parameters are often debatable. In general, there is hardly a choice of the tuning parameter that works across different data sets. If a data-driven method, such as cross-validation, is employed to select tuning parameters, it amounts to throwing the baby out with the bathwater: prediction will be subject to the selection effect. Secondly, predictions made using non-parametric models are usually only asymptotically valid (not finite-sample valid). Since all real-world applications are done in finitely many steps, any asymptotically correct model still leaves much to be desired regarding predictive validity.
@@
Given the above observations, it is natural to ask whether we can develop a predictive method with the following desiderata:
@@
The answer is affirmative. Indeed, conformal prediction (e.g. Vovk et al. 2005; Shafter and Vovk 2008) is a general machine learning strategy that can be used to create such a predictive method. Hong and Martin (2021) and Hong (2023) introduce conformal predction to insurance for the first time. These two papers develop predictive methods with all the above desiderata for unsupervised learning. The main goal of this paper is to develop an analogous method for the regression problem, i.e., a model-free and tuning-parameter-free method for making finite-sample valid predictions of a continuous response variable in supervised learning.
@@
The remainder of the paper proceeds as follows. Section 2 provides readers with the necessary background by reviewing conformal prediction. Then, Section 3 describes the proposed method in detail. Next, Section 4 uses numerical examples to demonstrate the excellent performance of the proposed method. Finally, Section 5 concludes the paper with concluding remarks. The Appendix contains some technical details.
@@ 2 Conformal prediction
Conformal prediction is a general machine learning strategy for generating prediction regions at a pre-determined coverage probability level. Shafer and Vovk (2008) give a user-friendly introduction to conformal prediction. For a comprehensive treatment, see Vovk et al. (2005). Conformal prediction can be used to tackle both supervised learning and unsupervised learning problems. Here we only review conformal prediction in supervised learning.
Suppose we observe pairs $Z_{i}=(X_{i},Y_{i})$, for $i=1,\ldots,n$, where $n$ is the sample size, $X_{i}\in\mathbb{R}^{p}$ is a vector of $p$ features, and $Y_{i}$ is the response variable. Throughout, $p$ is a positive integer. For any $n\geq 1$, we assume the finite sequence $Z_{1},Z_{2},\ldots,Z_{n}$ is exchangeable, i.e., the joint distribution of $(Z_{1},\ldots,Z_{n})$ is the same as the joint distribution of $(Z_{\sigma(1)},\ldots,Z_{\sigma(n)})$, where $\sigma:\{1,2,\ldots,n\}\rightarrow\{1,2,\ldots,n\}$ is any permutation on the set $\{1,2,\ldots,n\}$. Note that the exchangeability assumption is weaker than the widely used iid (independent identically distributed) assumption. In particular, if $Z_{1},\ldots,Z_{n}$ are iid, they are exchangeable (e.g., Chapter 1, Schervish 1995). Throughout, upper-case letters will denote random variables/vectors while the corresponding lower-case letters stand for their realizations.
Our chief goal is to predict the next response $Y_{n+1}$ at a randomly sampled feature $X_{n+1}$, based on past observations of $Z^{n}=\{Z_{1},\ldots,Z_{n}\}$. To apply conformal prediction, the actuary starts with a real-valued deterministic mapping $M$ of two arguments $(B,z)\mapsto M(B,z)$, where the first argument $B$ is a bag, i.e., a collection, of observed data and the second argument $z$ is a provisional value of a future observation to be predicted based on the data in $B$. $M$ is called a non-conformity measure. It measures the degree of non-conformity of the provisional value $z$ with the data in $B$. When the provisional value $z$ agrees with the data in $B$, $M(B,z)$ will be relatively small; otherwise, it will tend to be large. For example, if $z$ is real-valued and $B=\{z_{1},\ldots,z_{n}\}$, then we can take $M(B,z)=|\hat{m}_{B}(x)-y|$, where $\hat{m}_{B}(\cdot)$ is an estimate of the conditional mean function, $\mathsf{E}(Y\mid X=x)$, based on the bag $B$. There is no unique choice of the non-conformity measure. In practice, the non-conformity is chosen at the discretion of the actuary. Once a non-conformity measure is chosen, the actuary runs the conformal prediction algorithm—Algorithm 1 to predict the value of the next label $Y_{n+1}$ at a randomly sampled feature $X_{n+1}=x_{n+1}$.
1 Initialize: data $z^{n}=\{z_{1},\ldots,z_{n}\}$ and $x_{n+1}$, non-conformity measure $M$, and a provisional value $y$ of $Y$;
@@ 5       Compute $\mathsf{pl}_{z^{n}}(x_{n+1},y)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
8Return $\mathsf{pl}_{z^{n}}(x_{n+1},y)$ for each possible $y$;
@@
In Algorithm 1, $1_{E}$ denotes the indicator function of an event $E$. The quantity $\mu_{i}$, called the $i$-th non-conformity score, assigns a numerical value to $z_{i}$ to indicate how much $z_{i}$ agrees with the data in the $i$-th augmented bag $\widetilde{B}_{i}=z^{n}\cup\{z_{n+1}\}\backslash\{z_{i}\}$, where $z_{i}$ itself is excluded to avoid biases as in leave-one-out cross-validation. The function $\mathsf{pl}_{z^{n}}$ is called the plausibility function. It outputs a value between 00 and $1$, based on all non-conformity scores. The output of $\mathsf{pl}_{z^{n}}$ shows how plausible $z$ is a value of $Z_{n+1}$, based on the available data $Z^{n}=z^{n}$ and $X_{n+1}=x_{n+1}$. Using the output of the plausibility function $\mathsf{pl}_{z^{n}}$, the actuary can construct a $100(1-\alpha)\%$ conformal prediction region
@@
See the Appendix. ∎
@@
3 Proposed method

### v2

# Conformal prediction of future insurance claims in the regression problem

@@ Abstract
In the current insurance literature, prediction of insurance claims in the regression problem is often performed with a statistical model. This model-based approach may potentially suffer from several drawbacks: (i) model misspecification, (ii) selection effect, and (iii) lack of finite-sample validity. This article addresses these three issues simultaneously by employing conformal prediction—a general machine learning strategy for valid predictions. The proposed method is both model-free and tuning-parameter-free. It also guarantees finite-sample validity at a pre-assigned coverage probability level. Examples, based on both simulated and real data, are provided to demonstrate the excellent performance of the proposed method and its applications in insurance, especially regarding meeting the solvency capital requirement of European insurance regulation, Solvency II.
@@ 1 Introduction
Prediction is one of the most important inferential tasks for actuaries since it forms the basis for many key aspects of an insurer’s business operations, such as premium calculation and reserves estimation. According to Shmueli (2010), there are two key goals in data science and statistics: to explain and to predict. However, these two goals often warrant different approaches. For example, as demonstrated in Shmueli (2010), a wrong model, under some conditions, can even beat the oracle model in prediction, but the same cannot be said for explanation; see, also, p.13 of Tukey (1962) for a similar but broader philosophy. This paper only concerns prediction. In the extant insurance literature, prediction is often performed using either a parametric model or a non-parametric model (e.g., Frees et al. 2014). In the parametric model approach, the actuary posits a model, applies model selection tools to choose the “best” model, trains the chosen model, and finally makes predictions; see, for example, Claeskens and Hjort (2008) and Part I of Frees (2010). While this parametric approach has been widely applied in insurance, it has several drawbacks. First, the posited model may be misspecified, leading to grossly misleading predictions (Hong and Martin 2020). This prompted some authors to consider robust models (e.g., Brazauskas and Kleefeld 2016). Second, scholars recently discovered that model selection might introduce serious biases in prediction; this adverse effect is called the selection effect (e.g., Berk et al. 2013; Kuchibhotla et al. 2022); see, for example, Hong and Martin (2018a, b) for some simple illustrations in insurance. Though several proposals have been advanced to alleviate the selection effect, a satisfying answer has yet to be found (Kuchibhotla et al. 2022). Third, prediction is uncertain and should be probabilistically calibrated to achieve finite-sample validity (Gneiting and Katzfuss 2014; Vovk 2024); unfortunately, no parametric model is known to achieve finite-sample validity. For these reasons, some actuaries opt for a non-parametric model for predicting insurance claims (e.g., Jeon and Kim 2013; Hong and Martin 2017; Richardson and Hartman 2018). While most non-parametric models avoid both the model misspecification risk and the selection effect, they can still fall short of our expectations for two reasons. In the first place, many of these models contain some tuning parameters, and the choices of these tuning parameters are often debatable. In general, there is hardly a choice of the tuning parameter that works across different data sets. If a data-driven method, such as cross-validation, is employed to select tuning parameters, it amounts to throwing the baby out with the bathwater: prediction will be subject to the selection effect. Secondly, predictions made using non-parametric models are usually only asymptotically valid (not finite-sample valid). Since all real-world applications are done in finite steps, any asymptotically correct method still leaves much to be desired regarding predictive validity.
@@
Given the above observations, it is natural to ask whether we can develop a prediction method with the following desiderata:
@@
The answer is affirmative. Indeed, conformal prediction (e.g.,Vovk et al. 2005; Shafter and Vovk 2008) is a general machine learning strategy that can be used to create such a predictive method. Hong and Martin (2021) introduce conformal prediction to insurance for the first time. Hong (2023) and Chen et al. (2024) apply conformal prediction to credibility estimation and cat bond pricing, respectively. While Hong and Martin (2021) and Hong (2023) develop predictive methods with all the above desiderata for unsupervised learning, their methods, if adapted directly to supervised learning, would ignore the information supplied by predictors in the regression problem. This is certainly not desirable. Chen et al. (2024) make important contributions to the cat bond literature by applying conformal prediction to predict cat bond prices. But their method does not guarantee finite-sample validity of predictions. The reason is that their method cannot be implemented for every possible value of the response value; see Section 4 for a discussion on why finite-sample validity is not automatic when an actuary applies conformal prediction. Therefore, a predictive method, with the above three desiderata, for a regression problem is yet to be developed. The main goal of this paper is to propose such a method.
@@
The remainder of the paper proceeds as follows. Section 2 reviews the model-based approach and the model-free approach in the regression setup and argues why a model-free approach is desirable in prediction. Section 3 provides readers with the necessary background by reviewing conformal prediction. Then, Section 3 describes the proposed method in detail. Next, Section 4 uses simulation to show the excellent performance of the proposed method. It also gives a real-data example to demonstrate possible applications of the proposed method in insurance. Finally, Section 5 concludes the paper with concluding remarks.
2 Model-based approaches versus model-free approaches for prediction
In supervised learning, the true data-generating mechanism can be described as follows:
$Y=g^{\star}(X)+\epsilon,\quad i=1,\ldots,n,$ (1)
where $Y$ is the response variable, $X$ is the vector of $p$ features/predictors, $g^{\star}$ is an unknown deterministic function, $\epsilon$ is a random error term such that $\epsilon\sim F^{\star}$, and $F^{\star}$ is an unknown distribution function for $Y-g^{\star}(X)$. Throughout this article, $p$ is a positive integer, and upper-case letters will denote random variables/vectors while the corresponding lower-case letters stand for their realizations. Suppose we have observed independent and identically distributed (iid) pairs $Z_{i}=(X_{i},Y_{i})$, for $i=1,\ldots,n$ from (1), where $n$ is the sample size. Our chief goal is to predict the next response $Y_{n+1}$ at a randomly sampled feature $X_{n+1}$, based on past observations of $Z^{n}=\{Z_{1},\ldots,Z_{n}\}$. Since we aim at finite-sample valid predictions (see (4) in Section 3) and no point predictor can achieve this, we will only consider interval prediction. In general, there are two different types of approaches for performing prediction: a model-based approach and a model-free approach.
In a model-based approach, an actuary first posits a statistical learning model:
=$\mathcal{M}=\{(\mathcal{G},\mathcal{F})\},$
where $\mathcal{G}$ is a family of deterministic functions and $\mathcal{F}$ is a family of distribution functions. For example, in the linear regression model, we take $\mathcal{G}$ to be $\{g(x_{1},\ldots,x_{p})=a_{1}x_{1}+\ldots+a_{p}x_{p}\mid(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}\}$ and $\mathcal{F}$ to be $\{{\sf N}(0,\sigma^{2})|\sigma^{2}\in(0,\infty)\}$, where ${\sf N}(\mu,\sigma^{2})$ denotes the normal distribution function with mean $\mu$ and variance $\sigma^{2}$. Next, the actuary uses data to obtain estimators $\widehat{g^{\star}}$ of $g^{\star}$ and $\widehat{F^{\star}}$ of $F^{\star}$ such that $\widehat{g^{\star}}\in\mathcal{G}$ and $\widehat{F^{\star}}\in\mathcal{F}$. Finally, the actuary use $\widehat{g^{\star}}$ and $\widehat{F^{\star}}$ to create a prediction interval
$(L(Z^{n}),U(Z^{n})),$
where the lower bound $L(Z^{n})$ can be $\infty$- ∞ and the upper bound $U(Z^{n})$ can be +$\infty$+ ∞. If $\mathcal{G}$ and $\mathcal{F}$ both can be characterized by finitely many parameters, the model $\mathcal{M}$ is said to be a parametric model. For example, a linear model is a parametric model. If either $\mathcal{G}$ or $\mathcal{F}$ cannot be characterized by finitely many parameters, we say $\mathcal{M}$ is a non-parametric model. For example, if we take $\{g(x_{1},\ldots,x_{p})=a_{1}x_{1}+\ldots+a_{p}x_{p}\mid(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}\}$ and =$\mathcal{F}=\{F\mid\text{$F$ is a distribution function supported on $\mathbb{R}$}\}$, then $\mathcal{M}$ will be a non-parametric model.
Two challenges come into consideration when we employ a model-based approach for prediction: one is model misspecification, and the other is the effect of selection. If $g^{\star}\in\mathcal{G}$ and $F^{\star}\in\mathcal{F}$, we say the model $\mathcal{M}$ is well-specified or correct; otherwise, we say $\mathcal{M}$ is misspecified or wrong. In some cases, it can be unquestionable that a model is wrong. For example, if $Y_{i}$’s are observed insurance claims, then a linear model is patently misspecified because the response variable in a linear model can take negative values. However, we can never be certain that a model is correct because neither $g^{\star}$ nor $F^{\star}$ is observable. Therefore, even if a model seems to be reasonable, there is still a danger of model misspecification. Hong and Martin (2020) demonstrate that model misspecification can lead to grossly misleading predictions. One way to mitigate model misspecification is to apply some robust techniques in the spirit of Brazaukas and Kleefeld (2016). Another option is to use a non-parametric model. Either way, there is another danger lurking behind the scenes. When actuaries apply a parametric model, they often employ a model selection tool, such as AIC or BIC, to choose the “best” model. In such a case, the model is essentially treated as random (since the model selection process is data-dependent), but the existing formulas for predictions usually assume the model is fixed (i.e., conditional on the fact that the model has been chosen). This discrepancy can give rise to serious biases in prediction—a phenomenon called the selection effect; see Hong et al. (2018b) for an illustration. Note that if cross-validation is used to choose tuning parameters in a non-parametric model (e.g., bandwidth in a kernel density model), the selection effect might still wreak havoc on predictions. Some non-parametric models, such as the Dirichlet mixture of lognormal model in Hong and Martin (2017) may circumvent the selection effect. How to overcome the selection effect in a parametric model is an important topic of research in data science and statistics. At the time of writing, no efficient method has been found; see Kuchibhotla et al. (2022) for a review of recent results on the effect of selection.
The above two challenges in a model-based approach stem from the fact that prediction is conducted via a model. Therefore, one obvious way to circumvent these two difficulties is to perform prediction in one stroke: prediction is made from data directly without specifying a model. Such an approach is called a model-free approach for prediction. Technically, a model-free method for prediction is a mapping that maps data $Z^{n}$ to a prediction interval $(L(Z^{n}),U(Z^{n}))$. Figure 1 illustrates the difference between a model-based approach and a model-free approach. There are two paths from “Data” to “Prediction”. One first goes from “Data” to “Model” and then from “Model” to “Prediction”; this path denotes the model-based approach. The other path goes from “Data” directly to “Prediction”; it stands for the model-free approach.
DataModelPrediction
Figure 1: Model-based approach versus model-free approach for prediction.
A distinguishing mark of a model-free approach is that the notion of a model never enters the scene. Many machine learning techniques, such as decision trees, K-nearest neighbors, and neural networks, share this feature. Some elementary estimators in non-parametric statistics, such as sample mean and sample variance, have the same feature too, though they are designed for estimation, not prediction. A model-free approach for prediction can be considered a prediction method in non-parametric statistics. However, as mentioned above, a non-parametric statistical model might still be subject to the adverse impact of the selection effect. Moreover, non-parametric predictive models (e.g., Jeon and Kim 2013, Hong and Martin 2017, and Richardson and Hartman 2018) can ensure asymptotically valid prediction at best, and none of them guarantees finite-sample validity. Finally, non-parametric prediction methods exist in the literature (e.g., Wilks 1941; Fligner and Wolfe 1976; Frey 2013, Hong and Martin 2021, Hong 2023). But these methods are all designed for unsupervised learning. Adopting them directly for prediction in the regression problem means we need to disregard the information provided by $X_{1},\ldots,X_{n}$, which is undesirable. Therefore, we need to seek a new method in supervised learning that is (i) model-free, (ii) tuning-parameter-free, and (iii) guarantees finite-sample validity.
@@ 3 Conformal prediction
Conformal prediction is a general learning strategy designed for generating prediction regions at a pre-determined coverage probability level. Shafer and Vovk (2008) give a user-friendly introduction to conformal prediction. For a comprehensive treatment, see Vovk et al. (2005). Conformal prediction can be used to tackle both supervised learning and unsupervised learning problems. Here we only review conformal prediction in supervised learning.
Conformal prediction can be applied when data is exchangeable. For any $n\geq 1$, we say the finite sequence $Z_{1},Z_{2},\ldots,Z_{n}$ is exchangeable, if the joint distribution of $(Z_{1},\ldots,Z_{n})$ is the same as the joint distribution of $(Z_{\sigma(1)},\ldots,Z_{\sigma(n)})$, where $\sigma:\{1,2,\ldots,n\}\rightarrow\{1,2,\ldots,n\}$ is any permutation on the set $\{1,2,\ldots,n\}$. Note that the exchangeability assumption is weaker than the widely used iid (independent identically distributed) assumption. In particular, if $Z_{1},\ldots,Z_{n}$ are iid, they are exchangeable (e.g., Chapter 1, Schervish 1995).
Our goal is to predict the next response $Y_{n+1}$ at a randomly sampled feature $X_{n+1}$, based on past observations of $Z^{n}=\{Z_{1},\ldots,Z_{n}\}$. To apply conformal prediction, the actuary starts with a real-valued deterministic mapping $M$ of two arguments $(B,z)\mapsto M(B,z)$, where the first argument $B$ is a bag, i.e., a collection, of observed data and the second argument $z$ is a provisional value of a future observation to be predicted based on the data in $B$. $M$ is called a non-conformity measure. It measures the degree of non-conformity of the provisional value $z$ with the data in $B$. When the provisional value $z$ agrees with the data in $B$, $M(B,z)$ will be relatively small; otherwise, it will tend to be large. For example, if $z$ is real-valued and $B=\{z_{1},\ldots,z_{n}\}$, then we can take $M(B,z)=|\hat{m}_{B}(x)-y|$, where $\hat{m}_{B}(\cdot)$ is an estimate of the conditional mean function, $\mathsf{E}(Y\mid X=x)$, based on the bag $B$. There is no unique choice of the non-conformity measure. In practice, the non-conformity is chosen at the discretion of the actuary. Once a non-conformity measure is chosen, the actuary runs the conformal prediction algorithm—Algorithm 1 to predict the value of the next label $Y_{n+1}$ at a randomly sampled feature $X_{n+1}=x_{n+1}$.
1 Initialize: data $z^{n}=\{z_{1},\ldots,z_{n}\}$ and $x_{n+1}$, non-conformity measure $M$, and a grid of $y$ value;
@@ 5       Compute $\mathsf{pl}_{z^{n}}(x_{n+1},y)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
8Return $\mathsf{pl}_{z^{n}}(x_{n+1},y)$ for each $y$ on the grid;
@@
In Algorithm 1, $1_{E}$ denotes the indicator function of an event $E$. That is,
$1_{E}(x)=\left\{\begin{array}[]{ll}1,&\hbox{if $x\in E$,}\\ 0,&\hbox{otherwise.}\\ \end{array}\right.$
The quantity $\mu_{i}$, called the $i$-th non-conformity score, assigns a numerical value to $z_{i}$ to indicate how much $z_{i}$ agrees with the data in the $i$-th augmented bag $\widetilde{B}_{i}=z^{n}\cup\{z_{n+1}\}\backslash\{z_{i}\}$, where $z_{i}$ itself is excluded to avoid biases as in leave-one-out cross-validation. The function $\mathsf{pl}_{z^{n}}$ is called the plausibility function. It outputs a value between 00 and $1$, based on all non-conformity scores. The output of $\mathsf{pl}_{z^{n}}$ shows how plausible $z$ is a value of $Z_{n+1}$, based on the available data $Z^{n}=z^{n}$ and $X_{n+1}=x_{n+1}$. Using the output of the plausibility function $\mathsf{pl}_{z^{n}}$, the actuary can construct a $100(1-\alpha)\%$ conformal prediction region
@@
The proof is similar to its counterpart in the unsupervised learning context (e.g., Hong and Martin 2021). In great detail, $\mu_{1},\ldots,\mu_{n},\mu_{n+1}$ are exchangeable because $Z_{1},Z_{2},\ldots$ are exchangeable. Hence, the rank of $\mu_{n+1}$ follows the discrete uniform distribution on the set $\{1,\ldots,n,n+1\}$. By its definition, the plausibility function $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ is proportional to the rank of $\mu_{n+1}$. Therefore, $\mathsf{pl}_{Z^{n}}(Z_{n+1})$ also follows the discrete uniform distribution on the set $\{1/(n+1),2/(n+1),\ldots,1\}$. For a given $10<\alpha<$ 1, if $(n+1)\alpha$ is an integer, then $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}=\alpha$. Otherwise, we will have $\mathsf{P}^{n+1}\{\mathsf{pl}_{n+1}(Z_{n+1})\leq t_{n}(\alpha)\}\leq t_{n}(\alpha)<\alpha$. Therefore, (3) holds. ∎
@@
4 Proposed method for insurance applications